The sugar market has a supply curve with formula: Ps = 2 + 0.1Qs, and demand cuve: Pd = 55-0.32Qd

The government imposes a price floor of 22.9

Approximately what is the new consumer surplus?

Group of answer choices

1558

1610

1426

1592

The sugar market has a supply curve with formula: Ps = 2 + 0.1Qs, and demand cuve: Pd = 55-0.32Qd

The government imposes a price floor of 22.9

Approximately what is the dead weight loss?

Group of answer choices

118

141

130

108

The sugar market has a supply curve with formula: Ps = 2 + 0.1Qs, and demand cuve: Pd = 55-0.32Qd

What is the Price elasticity of demand (absolute value) using the average method, moving between P= 20 and P=21?

1.2

0.8

0.6

1.4

The sugar market has a supply curve with formula: Ps = 2 + 0.1Qs, and demand cuve: Pd = 55-0.32Qd

Between P= 20 and P=21 the Price elasticity of demand (absolute value) using the average method is

Group of answer choices

inelastic

elastic

The sugar market has a supply curve with formula: Ps = 2 + 0.1Qs, and demand cuve: Pd = 55-0.32Qd

What is the Price elasticity of supply using the average method, moving between P= 20 and P=21?

Group of answer choices

0.7

1.3

0.4

1.1

The sugar market has a supply curve with formula: Ps = 2 + 0.1Qs, and demand cuve: Pd = 55-0.32Qd

Between P= 20 and P=21 the Price elasticity of supply using the average method is

Group of answer choices

inelastic

elastic

The sugar market has a supply curve with formula: Ps = 2 + 0.1Qs, and demand cuve: Pd = 55-0.32Qd

given a market price around p=20c/pound, a tax imposed on the supplier will mostly be paid by

Group of answer choices

both consumer and supplier equally

the consumer

the supplier
Answer #1

At P = 22.9, Quantity sold is (55 – 22.9)/0.32 = 100.3125. Hence CS = 0.5*(55 – 22.9)*100.3125 = 1610.

Old quantity is 126.1905 units and difference between price buyers willing to pay and price sellers willing to receive at 100.3125 is (22.9 – 12.03125) = 10.86875. DWL = 0.5*(126.1905 – 100.3125)*10.86875 = 141

At P = 20, Qd = 109.375 and at P = 21, Qd = 106.25. This gives ed = -0.59 of 0.6.

Inelastic since ed < 1

At P = 20, Qs = 180 and at P = 21, Qs = 190. This gives es = 1.1

Elastic since es > 1

Consumers